Part A) 1.23g of Ba(OH)2 is dissolved in enough pure water at 25 degrees C to make a 50.0mL solution. What is the pH of the solution?
Part B) the 50.0mL solution of Ba(OH)2 in Part A is titrated with a 0.287mol/L solution of hydrochloric acid (HCl). what is the pH after the addition of 45.0mL of HCl?
1
Expert's answer
2020-11-02T08:55:52-0500

A) V = 50 mL = 0.05 L

M(Ba(OH)2) = 171.34 g/mol

m(Ba(OH)2) = 1.23 g

n = m/M

n(Ba(OH)2) "= \frac{1.23}{171.34} = 0.007178 \;mol"

[Ba(OH)2)] "= \frac{0.007178 \;mol}{0.05 \;L} = 0.143 \;M"

Ba(OH)2 → Ba2+ + 2OH-

Ba(OH)2 is a strong base

[OH-] = 2[Ba(OH)2)]

[OH-] = "2 \times 0.143 \;M = 0.286 \;M"

[H+] "= \frac{1.0 \times 10^{-14}}{[OH-]}" at 25 degrees C

[H+] "= \frac{1.0 \times 10^{-14}}{[0.286]} = 3.4965 \times 10^{-14} \;M"

pH = -log[H+]

pH = "-log(3.4965 \times 10^{-14})"

pH = 13.45

B) [HCl] = 0.287 mol/L

V(HCl) = 45 mL = 0.045 L

n(HCl) "= 0.287 \;mol \times 0.045 \;L = 0.012915 \;mol"

[Ba(OH)2)] = 0.143 M

V(Ba(OH)2) = 50 mL = 0.05 L

n(Ba(OH)2) "= 0.143 \; mol/L \times 0.05 \;L = 0.00715 \;mol"

n(OH-) = 2n(Ba(OH)2)) "= 2 \times 0.00715 \;mol = 0.0143 \;mol"

n(H+) = 0.012915 mol

Remaining n(OH-) = 0.0143 – 0.012915 = 0.001385 mol

Total V = 0.05 L + 0.45 L = 0.095 L

[OH-] "= \frac{0.001385 \;mol}{0.095 \;L} = 0.01457 \; M"

[H+] "= \frac{1.0 \times 10^{-14}}{[OH-]}"

[H+] "= \frac{1.0 \times 10^{-14}}{[0.01457]} = 68.6341 \times 10^{-14} \;M"

pH = -log[H+]

pH "= -log(68.6341 \times 10^{-14})"

pH = 12.16